OpenStudy (anonymous):

The functions f, g ,h and are defined as follows. f(x)=5+ sqrt x-3 g(x)=|-2/3x-15| h(x)= 4+x^2/x+4 Find f(5), g(9), and h(2).

OpenStudy (anonymous):

simply plug in the value everywhere there is x. f(5) = 5 + sqrt 5 - 3

OpenStudy (anonymous):

so the answer would be: f(5)=5+cuberoot x-3? Am I doing this right?

OpenStudy (mertsj):

\[f(x)=5+\sqrt{x-3}\]

OpenStudy (mertsj):

Is that right?

OpenStudy (anonymous):

Yes, Mertsj, thats what the original problem looks like.

OpenStudy (anonymous):

I have g(9)=21 for g if i am doing this right?

OpenStudy (mertsj):

Then to find the f(5) replace x with 5

OpenStudy (mertsj):

That is correct for the g(9)

OpenStudy (anonymous):

\[f(x)=5+ \sqrt{x-3}\]\[g(x)={{-2}\over{3x-15}}\]\[h(x)={{4+x^2}\over{x+4}}\] to find \(f(5)\), \(g(9)\) and \(h(2)\) simply plug in 5, 9, and 2 for \(x\) in the functions: \[f(5)=5+ \sqrt{5-3}\]\[\color{purple}{f(5)=5+ \sqrt{2}}\] \[g(9)={{-2}\over{3(9)-15}}\]\[g(9)={{-2}\over{27-15}}\]\[g(9)={{-2}\over{12}}\]\[\color{purple}{g(x)=-{{1}\over{6}}}\] \[h(2)={{4+2^2}\over{2+4}}\]\[h(2)={{4+4}\over{6}}\]\[h(2)={{8}\over6}\]\[\color{purple}{h(2)={{4}\over3}}\] hope this helps! :)

OpenStudy (anonymous):

oh wait..i didnt see the absolute value signs for g(9)...ignore that one

OpenStudy (anonymous):

thank you all!